OpenStudy (anonymous):

what is the solution of the equation 2^x+1

OpenStudy (anonymous):

2^x+1 is not an equation

OpenStudy (anonymous):

it is a function it will have lots of answers depending on what x you put into it this one is just an arc with a horizontal asymptote at 1 (because of the +1); y intercept at 2 because 2^0 = 1 +1 = 2 and then an exponential trip to oo after that